OpenStudy (anonymous):

f(x)=2x+5; g(x)=9x-7 (1.)(f-g)(x)=? (2.)what is the domain of f-g? a.{x|x>0}(>or equal to) b. {x|x>-5/2 c. {x|x is any real number} d. {x|x= -5/2}(not equal to) 3.what is the domain of f/g? a.{x|x=7/9}(not equal to sign) b.{x|x>7/9} c.{x|x>0)(>or equal to) d.{x|x is any real number}

OpenStudy (anonymous):

Do you mean (f(x)-g(x))*x?

OpenStudy (anonymous):

If so, you get: (-7x+12)*X=-7x^2+12x

OpenStudy (anonymous):

the top is the question and then you have to answer the bottom 3

OpenStudy (anonymous):

I see, what you want to do is use f-g as a reference for the domains etc. f-g=12-7x

OpenStudy (anonymous):

As for (2.) I'm fairly sure the answer is c. (\[x \in \mathbb{R}\]

OpenStudy (anonymous):

f/g = (2x+5)/(9x-7) Here you're more limited as 9x-7 cant be 0, unless you want a great explosion from dividing by zero! :D

OpenStudy (anonymous):

9x-7=0 9x=7 x=7/9 Therefore x can't be 7/9. The answer would be A.

OpenStudy (anonymous):

awesome thank you soo much!!

OpenStudy (anonymous):

No problem! Just make sure you understand my reasoning - if you don't, ask!